OpenStudy (anonymous):

write a third-degree polynominal using these numbers 1, 2-i

OpenStudy (anonymous):

if the polynomial has only real coeffecients then roots would be x= 1, 2 - i, 2+i p(x) = (x-1)(x -2+i)(x - 2-i)

OpenStudy (anonymous):

how do you write that using exponents??

OpenStudy (anonymous):

you multiply = > (x-1) * (x^2 + 5 -4x) = (x^3 +5x - 4x^2 - x^2 -5 +4x) = x^3 -5x^2 +9x - 5

OpenStudy (anonymous):

thank you